OpenStudy (gavin39):

Need a quick answer What is the simplified form of 4x plus 2 over x plus 5 plus the fraction 3x minus 1 over x plus 5 ? x plus 1 over x plus 5 x plus 3 over x plus 5 7x plus 1 over x plus 5 7x plus 3 over x plus 5

OpenStudy (gavin39):

Will gie medal

OpenStudy (agent0smith):

Use equations not words.

OpenStudy (studygurl14):

\[\frac{ 4x + 2 }{ x + 5 } + \frac{ 3x - 1 }{ x + 5 }\] \[\frac{ (4x + 2) + (3x - 1) }{ x + 5 }\] \[\frac{ 7x + 1 }{ x + 5 }\]

OpenStudy (gavin39):

@StudyGurl14 could you help me with a few more

OpenStudy (studygurl14):

sure @gavin39